# ComplexityMeasures.jl: scalable software to unify and accelerate entropy and complexity timeseries analysis

### v1

@@ 5 Conclusions
In conlusion in this paper we introduced ComplexityMeasures.jl. We highlighted the software’s scalability, both in terms of being able to compute countless complexity measures with little source code, or almost no effort from the user, but also in terms of computational performance, being up to 1,000x faster than alternatives. The performance aspect enables practical applications that were previously limited by computational capacity, such as extensive null hypothesis testing using surrogate data on large time series ensembles. We also highlighted its composable design that allows creating new types of complexity measures “on the fly”, and its extensibility, which allows any new measure or estimator to be instantly usable within the wider ecosystem. For this, we believe that wider adoption of ComplexityMeasures.jl can be an invaluable asset for both academics and industry, wherever estimating complexity measures is useful.
@@
We also wish to highlight the openness in the development of ComplexityMeasures.jl. ComplexityMeasures.jl was purposefully built from the ground up on Github as a community effort. After we, the authors, crossed path in various open source projects trying to implement similar functionality, we quickly realized that there were many commonalities between our disparate research fields. We also realized that the most efficient way forward to ensure solid, reproducible research, and to achieve a good overview of the field, was to establish a common software framework for complexity quantification. Since its conception more than three years ago, ComplexityMeasures.jl’s Github repository has seen over 200 pull requests and over 140 issues have been closed. We follow best open source practices, and greatly value openness and community contributions — principles which we share with the software consortium for scientific computing developed by the JuliaDynamics organization.
@@
It is our hope that in the future, researchers that develop new complexity measures contribute them directly to ComplexityMeasures.jl when submitting their paper for review. This has numerous benefits for both the researchers themselves, and for the wider community. For the researchers, their new method is instantly accessible to an established pool of users, and citable via BiBTeX integration, generating the contributing researchers citations and recognition. The method is also instantly integrated with a wide range of estimators, allowing further research and study. For the community, the method becomes part of a well-tested, well-documented, and established software, enhancing reliability and accessibility. The integration of ComplexityMeasures.jl with a wider ecosystem such as DynamicalSystems.jl also allows the community to easily test whether the claims of a new paper are accurate and robust w.r.t. variability in the parameters or input data (e.g., see the applications of Sec.3.2 and 3.3). Lastly, this approach also promotes openness in academic code, an aspect that we believe is important in its own right, beyond complexity measures.
@@
using ComplexityMeasures # at least version 3.5
@@
In Table 2 we list the content of ComplexityMeasures.jl in terms of available outcome spaces, information and complexity measures, and estimators thereof. In this subsection we use this table to count how many total complexity measures can be estimated with ComplexityMeasures.jl, version 3.5.0. We advise the reader to visit the software documentation for always-up-to-date measure counts estimates that are computed fully programmatically from the software source code.
@@
-2.25in0in Abstract type Concrete implementations OutcomeSpace Uni- or multivariate timeseries: UniqueElements, ValueBinning, OrdinalPatterns, WeightedOrdinalPatterns, AmplitudeAwareOrdinalPatterns, Dispersion, CosineSimilarityBinning, BubbleSortSwaps, SequentialPairDistances, TransferOperator, NaiveKernel, WaveletOverlap, PowerSpectrum; Spatiotemporal timeseries: SpatialDispersion, SpatialOrdinalPatterns, SpatialBubbleSortSwaps ProbabilitiesEstimator RelativeAmount, Shrinkage, BayesianRegularization, AddConstant InformationMeasure Shannon, Renyi, Tsallis, Curado, Kaniadakis, StretchedExponential, ShannonExtropy, RenyiExtropy, TsallisExtropy DiscreteInfoEstimator Generic: PlugIn, Jackknife; Shannon-entropy specific: MillerMadow,HorvitzThompson, Schuermann, GeneralizedSchuermann, ChaoShen DifferentialInfoEstimator Shannon entropy: KozachenkoLeonenko, Kraskov, Goria, Gao, Zhu, ZhuSingh, Lord, AlizadehArghami, Correa, Vasicek, Ebrahimi; Shannon, Rényi or Tsallis entropy: LeonenkoProzantoSavani ComplexityEstimator ApproximateEntropy, SampleEntropy, LempelZiv76, MissingDispersionPatterns, StatisticalComplexity, ReverseDispersion, BubbleEntropy
Table 2: Central abstract types in ComplexityMeasures.jl and their concrete implementations for software version 3.5.0.
@@
There are 11 discrete information measure definitions in the software. Every discrete information measure can be estimated with either of the two generic estimators JackKnife or PlugIn. For Shannon entropy, five additional dedicated estimators are implemented. This means that in total we have $11\cdot 2+5=27$ ways to estimate a discrete information measure from a PMF. Since all discrete information measures are functionals of PMFs, the number of ways to estimate a discrete information measure from data is $27\cdot 46=1242$.
@@
Finally, the software provides 7 complexity estimators that are not functionals of PMFs. One of them, StatisticalComplexity, must be handled separately, because in our software, we made the innovation to make the statistical complexity computable from any configuration of input outcome spaces, probabilities estimators, discrete information measure definition and estimator. To be a bit conservative with counting, we here only count the variability of the statistical complexity arising from varying outcome spaces and discrete information measure definitions, as these two would have the strongest impact on the statistic. This gives $16\cdot 11=176$ variants of statistical complexity, and the total number of non-probability-based complexity measures is thus $176+6=182$.
@@
This makes the grand total of measures that one can estimate with ComplexityMeasures.jl equal to: $92+182+14+1242=1530$.
@@ Acknowledgments
We would like to acknowledge all volunteer contributors to ComplexityMeasures.jl, which can be found on its GitHub page under the URL: https://github.com/JuliaDynamics/ComplexityMeasures.jl/graphs/contributors. KAH would like to thank the Earth System Evolution group at the University of Bergen, in particular Bjarte Hannisdal and David Diego, for creating a thriving environment where scientific software development becomes a natural part of the research process.
@@
The figures and the performance numbers quoted in the comparison table are fully reproducible. The codebase that produced them can be found in [62].

### v3

@@ 5 Conclusions
In this paper, we introduced ComplexityMeasures.jl, the reasoning behind its design, and its efficacy with respect to alternatives. The advantages and unique features of ComplexityMeasures.jl can be summarized as follows:
  • Modular and composable design based on the mathematically rigorous formulation of a complexity measure. This design allows computing complexity measures that have not even been published yet in the literature.
  • Easily scalable in terms of implementing new measures devised in new research without writing much new code.
  • Intelligent utilization of Julia’s multiple dispatch system [40] to create an incredibly lean codebase averaging only 2.3 lines of code per potential measure to be estimated.
  • Exceptional computational performance, up to 1,000x faster than alternatives. This enables practical applications that were previously limited by computational capacity, such as extensive null hypothesis testing using surrogate data on large timeseries ensembles.
  • Extensively tested through automated testing (continuous integration), providing a trustworthy environment.
Due to these features, we believe that a wider adoption of ComplexityMeasures.jl can be an invaluable asset for both academics and industry, wherever estimating complexity measures is useful.
@@
We also wish to highlight the openness in the development of ComplexityMeasures.jl. ComplexityMeasures.jl was purposefully built from the ground up on Github as a community effort. After we (the authors) crossed paths in various open source projects trying to implement similar functionality, we quickly realized that there were many commonalities between our disparate research fields. We also realized that the most efficient way forward to ensure solid, reproducible research, and to achieve a good overview of the field, was to establish a common software framework for complexity quantification. Since its conception more than four years ago, ComplexityMeasures.jl’s Github repository has seen over 200 pull requests and over 140 issues (public discussions) have been addressed, implemented, and closed. We follow best open-source practices and greatly value openness and community contributions (see Acknowledgments).
@@
It is our hope that, in the future, researchers that develop new complexity measures contribute them directly to ComplexityMeasures.jl when submitting their paper for review. This has numerous benefits for both the researchers themselves, and for the wider community. For the researchers, their new method is instantly accessible to an established pool of users, and citable via BiBTeX integration, generating the contributing researchers citations and recognition. The method is also instantly integrated with a wide range of estimators, allowing further research and study. For the community, the method becomes part of a well-tested, well-documented, and established software, enhancing reliability and accessibility. The integration of ComplexityMeasures.jl with a wider ecosystem such as DynamicalSystems.jl also allows the community to easily test whether the claims of a new paper are accurate and robust w.r.t. variability in the parameters or input data (e.g., see the applications of Sec.3.2 and 3.3). Lastly, this approach also promotes openness in academic code, an aspect that we believe is important in its own right, beyond complexity measures.
@@
using ComplexityMeasures # at least version 3.7
@@
In Table 2 we list the content of ComplexityMeasures.jl in terms of available outcome spaces, information and complexity measures, and estimators thereof. In this subsection we use this table to count how many total complexity measures can be estimated with ComplexityMeasures.jl, version 3.7.0. We advise the reader to visit the software documentation for always-up-to-date measure counts estimates that are computed fully programmatically from the software source code [46].
@@
-2.25in0in Abstract type Concrete implementations OutcomeSpace Uni- or multivariate timeseries: UniqueElements, ValueBinning, OrdinalPatterns, WeightedOrdinalPatterns, AmplitudeAwareOrdinalPatterns, Dispersion, CosineSimilarityBinning, BubbleSortSwaps, SequentialPairDistances, TransferOperator, NaiveKernel, WaveletOverlap, PowerSpectrum; Spatiotemporal timeseries: SpatialDispersion, SpatialOrdinalPatterns, SpatialBubbleSortSwaps ProbabilitiesEstimator RelativeAmount, Shrinkage, BayesianRegularization, AddConstant InformationMeasure Shannon, Renyi, Tsallis, Curado, Kaniadakis, StretchedExponential, ShannonExtropy, RenyiExtropy, TsallisExtropy, FluctuationComplexity DiscreteInfoEstimator Generic: PlugIn, Jackknife; Shannon-entropy specific: MillerMadow,HorvitzThompson, Schuermann, GeneralizedSchuermann, ChaoShen DifferentialInfoEstimator Shannon entropy: KozachenkoLeonenko, Kraskov, Goria, Gao, Zhu, ZhuSingh, Lord, AlizadehArghami, Correa, Vasicek, Ebrahimi; Shannon, Rényi or Tsallis entropy: LeonenkoProzantoSavani ComplexityEstimator ApproximateEntropy, SampleEntropy, LempelZiv76, MissingDispersionPatterns, StatisticalComplexity, ReverseDispersion, BubbleEntropy
Table 2: Central abstract types in ComplexityMeasures.jl and their concrete implementations for software version 3.7.
@@
There are 12 discrete information measure definitions in the software. Every discrete information measure can be estimated with either of the two generic estimators JackKnife or PlugIn. For Shannon entropy, five additional dedicated estimators are implemented. This means that in total we have $12\cdot 2+5=29$ ways to estimate a discrete information measure from a PMF. Since all discrete information measures are functionals of PMFs, the number of ways to estimate a discrete information measure from data is $29\cdot 46=1334$.
@@
Finally, the software provides 7 complexity estimators that are not functionals of PMFs. One of them, StatisticalComplexity, must be handled separately, because in our software, we made the innovation to make the statistical complexity computable from any configuration of input outcome spaces, probabilities estimators, discrete information measure definition and estimator. To be a bit conservative with counting, we here only count the variability of the statistical complexity arising from varying outcome spaces and discrete information measure definitions, as these two would have the strongest impact on the statistic. This gives $16\cdot 12=192$ variants of statistical complexity, and the total number of non-probability-based complexity measures is thus $192+6=198$.
@@
This makes the grand total of measures that one can estimate with ComplexityMeasures.jl equal to: $92+198+14+1334=1638$. This total is also estimated programmatically in the online documentation of ComplexityMeasures.jl (so see the online documentation for an up-to-date estimation [46]).
@@ Acknowledgments
We would like to acknowledge all volunteer contributors to ComplexityMeasures.jl which can be found on its GitHub page under the URL: https://github.com/JuliaDynamics/ComplexityMeasures.jl/graphs/contributors. GD acknowledges funding from UKRI’s Engineering and Physical Sciences Research Council, grant no. EP/Y01653X/1 (grant agreement for an EU Marie Sklodowska-Curie Postdoctoral Fellowship). KAH would like to thank the Earth System Evolution group at the University of Bergen, in particular Bjarte Hannisdal and David Diego, for creating a thriving environment where scientific software development becomes a natural part of the research process. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@
The figures and the performance numbers quoted in the comparison table are fully reproducible. The codebase that produced them can be found in [66]. Data collection and analysis method complied with the terms and conditions of the source of data.